OpenStudy (anonymous):

Solve 4x + 7 = 6x – 1 .. how do you solve this ? "/

OpenStudy (anonymous):

-2x=-8 x=(-8)/(-2)=4

OpenStudy (anonymous):

Solve 4 ^ x + 7 = 6 ^ x – 1 <- this is how it is written , the answer choices are x ≈ 0.04 x ≈ 0.05 x ≈ 19.51 x ≈ 28.35

OpenStudy (anonymous):

I calculated that x = 1.58

OpenStudy (anonymous):

\[4^{x}+7=6^{x}-1\] is this how it was written?